# Cybersecurity in Critical Infrastructures: A Post-Quantum Cryptography Perspective

### v1

@@ III.1.2 Asymmetric Key Schemes
  • Encryption Algorithm (Enc): Bob with the public key encrypts the plaintext (the message) generating the ciphertext (Ct). He then proceeds to publically communicate the protected message:
@@
This is the general scheme of a asymmetric key cryptosystem, the security of this scheme relies on the assumption that Eve is not able to recover the original plaintext by just using the public key and the ciphertext. Such thing relies on a mathematical problem that is hard to solve but for which it is easy to prove if a given solution is correct or not. Those are usually referred as one-way functions. They are designed in such a way that Eve can not decipher the message in reasonable time without the private key and she can not get any information about the private key by means of the public information. The most famous, and widely used, scheme is the RSA cryptosystem, developed by R.L. Rivest, A. Shamir and L. Adleman in 1979 [6]. Its security is based on the factorization of large numbers in their prime factors. Other used schemes are based on ECC such as the Elliptic Curve Diffie-Hellman (ECDH) scheme [45].
III.1.3 Security Notion
Proving the security of a cybersecurity scheme is not a trivial problem, i.e. mathematically proving that a function is indeed an One-Way function is no a simple task. In fact, one of the “Millenium Prize Problems”, known as the $P\overset{?}{=}NP$ problem [46], involves to proof if the set of the problems whose solution is hard to find is the same to the set of problems whose solution is easy to check. However, the security of cryptosystems can be studied by means of some security notions. There are three cases which have to be taken into account to study a cryptosystem’s security. In all of them, the adversary who wants to break the security is modeled by a probabilistic polynomial time algorithm. The scheme is secure if the algorithm has no advantage for discovering the secret over a random guesser algorithm, i.e. it is said Eve has a negligible advantage if she guess the correct answer with probability $1/2+\epsilon(k)$, where $k$ is a security parameter and $\epsilon(k)$ is a negligible function. Following this logic, the three security notions considered are:
@@ III.2 Quantum Apocalypses
As explained before, the key exchange algorithm for establishing the secret key of a symmetric protocol is done with a public key (asymmetric) cryptosystem. In this sense, doubling the size of the key to be shared increments the complexity of the key generation algorithm of the public key scheme. However, increasing the computational cost is feasible in order to hold the security in almost all the cases and, hence, this is not problematic. However, the security of a public key cryptosystem relies on the complexity of the hard problem which has to be solved to get the private key. For example, a brute force attack to RSA consists in trying all primes $p$ and checking whether $p$ is a factor of $N$, requiring $\sqrt{N}$ attempts in the worst case scenario, which is exponential in the digits of $N$ ($d$) and, thus, is an unfeasible task for any classical computer. The most efficient classical algorithm to solve such problem, known as the general number field sieve [51], achieves a complexity of $\mathcal{O}\left(e^{\sqrt{ln(n)ln(ln(n))}}\right)$ asymptotically [52], using this method up to RSA-250 (250 digits) have been factorized satisfactorily [53].
@@
A large enough fault-tolerant quantum computer could run Shor’s algorithm, which is able to factor a number in its primes making advantage of the laws of quantum mechanics. The first large enough quantum computer will be able to solve factorization problem using $10d$ logical qubits [^3] Note that this number refers to logical or “noiseless” qubits [54]. For quantum computers to work, quantum error correction is required implying that many more physical or “noisy” qubits will be required for an implementation of the Shor algorithm that cracks RSA [55]., where $d$ is the number of digits, with a complexity of $\mathcal{O}(d^{3})$ [52]. A comparison between both algorithms, in terms of their bounds in number of steps, is represented in FIG. 4, where the huge difference (exponential speedup) between both algorithms when the number of digits increases can be easily observed.
Therefore, Shor’s algorithm is a potential threat to the security of all communications over the world, and will be a real problem in the near future since it is able to break most used public key cryptosystems, RSA and ECC schemes. As a result, governments and private companies, such as technological giants IBM, Google and Microsoft, to cite some of them; are dedicating a humongous amount of economical and human resources to develop a large enough functional quantum computer (fault-tolerant) able to compute Shor’s algorithm and break the current cybersecurity. Hence, new cryptography schemes are required in order to protect our communications and documents in the post-NISQ [^4] Noisy intermediate-scale quantum era: refers to time scale in which quantum advantage has been firmly proven, but the quantum computers available are still small and too noisy to offer the full potential of quantum computing [56]. NISQ is where we stand right now. quantum era. Another issue to take into account is the fact that some people speculate about the possibility that some global agents are storing encrypted communications for decrypting it in the future once fault-tolerant quantum computers are available. Thus, finding an alternative to RSA and ECC cryptography resistant to quantum attacks is a very relevant problem nowadays. Moreover, it is estimated that a secret holds its value for 15 years [17] and, thus, it is necessary to consider public key cryptosystems based on hard problems that can not be solved efficiently by quantum computers as soon as possible, i.e. it is necessary to implement quantum resistant algorithms 15 years before the first functional quantum computer is available. Google, IBM, China and Xanadu have shown quantum advantage using their computers and they expect a huge increase in their quantum computer capability during the following years [10, 11, 12, 13, 14]. To sum up, we are at a critical moment for cryptography.
@@ IV Post-Quantum Cryptography
Post Quantum Cryptography (PQC) refers to classical cryptographic methods based on hard problems whose solution cannot be found in polynomial time by neither classical nor quantum computers. The hardness of the problem is defined by the computational complexity of an algorithm capable to solving it. In this sense, a quantum computer should not provide any advantage (or such should be almost negligible) in solving the hard problem that stands at the core of a PQC protocol.
@@
IV.1 Hash-based cryptography
@@
Hash functions are usually employed to create digital signatures, deemed as hash-based signatures (HSB), providing authenticity and integrity to the communication. The first signature scheme using hash-functions was introduced by Lamport in 1979 [60]. Those functions can be classified into One Way Hash Functions (OWHF), Collision Resistant Hash Functions (CRHF) and Universal One way Hash Functions (UOWHF). The general scheme of a hash function cryptosystem is composed by three algorithms, Gen, Enc and Dec in the case of encryption or Gen, Sign and Vry in the case of a digital signature scheme:
@@ IV.1.1 Security Notion
IV.1.2 PQC protocols
There are two types of hash-based digital signatures (HBS): the One-Time Signature (OTS) scheme and the Multi-Time Signature (MTS) scheme, sometimes called statefull HBS and stateless HBS, respectively. The principal difference between them is the times that a secret key can be used without losing the security assumption. Only Europe, Japan and the United States have a HBS algorithm in their standardization process, SPHINCS+ [62], which is a MTS signature scheme.
The parameters of such algorithm are available in Table 4.
IV.2 Lattice-based cryptography
@@
Another lattice-based cryptosystem is based on the Learning With Errors (LWE) problem, it is defined below, which was proposed by Regev [65]. The three algorithms involving such cryptography scheme are:
@@ IV.2.1 Security Notion
Despite of the existence of an algorithm which finds the non-zero vector of the SVP in time $2^{\mathcal{O}(n)}$ [66], there is no quantum algorithm providing an exponential speedup. Indeed, it is worth noting that the SVP problem can be reduced to the CVP problem. In 1996, Ajtai et al. showed that there is a connection between the worst-case problems and the average-case problems in lattice-based cryptography and explain how to construct hard lattice instances from random instances [67]. This connection implies that if there is a probabilistic algorithm that solves the hard problems in the average-case, then there exists a solution for the worst-case scenario. However, it has been proven that there is no probabilistic algorithm in polynomial time for the worst-case scenario. Finally, they conclude that a lattice-based cryptography scheme based on average-case problems can be designed with the security of the worst-case problems. The average-case problems related with lattices are:
@@
IV.2.2 PQC protocols
Lattice-based cryptography can be splitted in two diffferent algorithms: NTRU, developed by mathematicians Jeffrey Hoffstein, Jill Pipher, and Joseph H. Silverman in 1996 [64]; and Learning With Errors (LWE) first introduced by Oded Regev in 2009 [65]. Lattice-based cryptography is the most promising quantum resistant cryptosystem, for encryption and signatures. This can be seen as different countries are proposing a large amount of lattice-based protocols to be be standardized. The NTRU encryption protocols that have been proposed to be standardized are: NTRU [69], NTRU-HRSS [70], NTRU-Prime [71] and NTRU+ [72]; and the following NTRU protocols have been also proposed to generate signatures: Falcon [73], FatSeal [74], Peregrine [75] and SOLMAE [76]. On the other hand, the LWE encryption protocols proposed to be standardized are: CRYSTALS-Kyber [27], Saber [77], FrodoKEM [78], LAC PKE [79], Aigis-Enc [80], AKCN-MLWE [81], TALE [82], AKCN-E8 [83], SCloud [84], SMAUG [85] and TiGER [86]; and the following LWE protocols have been also proposed to generate signatures: Crystal-Dilithium [87], Aigis-Sig [88], Mulan [89], GCKSign [90], NCC-Sign [91] and HAETAE [92].
The parameters of those algorithms are available in Table 3.
IV.3 Code-based cryptography
@@ IV.3.1 Security Notion
IV.3.2 PQC protocols
Two Niederreiter protocols have been proposed to be standardized: BIKE [100] proposed to the NIST standardization process [101] and PALOMA [102] proposed to the Korea PQC standardization process [103]. In the case of McEllice protocols, NIST process is considering Classical McEllice [104] and HQC protocols [105], and South Korea is considering [106] and [107]. Enhanced pqsigRm [108] is also being evaluated as a code-based signature scheme in Korean PQC standardization process.
The parameters of those algorithms are available in Table 5.
IV.4 Multivariate cryptography
@@ IV.4.1 Security Notion
IV.4.2 PQC protocols
There are two multivariate cryptosystems proposed to be standardized: Rainbow [112] and GeMSS [113]. Both of them have been proposed in the NIST standardization process. On the other hand, South Korea has proposed a multivariate crypyptography protocol called MQ-sign [114].
The parameters of those algorithms are available in Table 6.
IV.5 Isogeny-based cryptography
@@ IV.5.1 Security Notion
The best algorithm to solve the SSIP and, hence, that breaks isogeny-based cryptography, is based on meet-in-the-middle attack and its complexity is $\mathcal{O}(p^{1/4})$, which requires $\mathcal{O}(p^{1/4})$ storage capability [121], or, in the case of SIKE, its security is based on van Oorschot-Wiener (vOW) golden collision finding algorithm [122]. The best quantum algorithm is able to find the secret with $\mathcal{O}(p^{1/6})$, which is not a notorious advantage and does not compromise isogeny-based cryptography [121].
IV.5.2 PQC protocols
Supersingular Isogeny Key Encapsulation (SIKE) [123] is an encryption scheme based on Supersingular Isogeny Diffie-Hellman protocol proposed for the NIST standardization process. On the other hand, FIBS is a signature scheme based on the same protocol proposed in the South-Korean standardization process [124].
IV.6 Multiparty Computation protocol and Graph-based cryptography

### v2

@@ III-A2 Asymmetric Key Schemes
  • Encryption Algorithm (Enc): Bob with the public key encrypts the plaintext (the message Msg) generating the ciphertext (Ct). He then proceeds to communicate publicly the protected message:
@@
This is the general scheme of an asymmetric key cryptosystem, the security of which relies on the assumption that Eve is not able to recover the original plaintext by just using the public key and the ciphertext. Such thing relies on a mathematical problem that is hard to solve but for which it is easy to prove if a given solution is correct or not. Those are usually referred as one-way functions. They are designed in such a way that Eve can not decipher the message in reasonable time without the private key and she can not get any information about the private key by means of the public information. The most famous, and widely used, scheme is the RSA cryptosystem, developed by R.L. Rivest, A. Shamir and L. Adleman in 1979 [5]. Its security is based on the factorization of large numbers in their prime factors. Other used schemes are based on ECC such as the Elliptic Curve Diffie-Hellman (ECDH) scheme [57].
III-A3 Security Notion
Proving the security of a cybersecurity scheme is not a trivial problem, i.e. mathematically proving that a function is indeed an One-Way function is not a simple task. In fact, one of the “Millenium Prize Problems”, known as the $P\overset{?}{=}NP$ problem [58], involves to prove if the set of the problems whose solution is hard to find is the same to the set of problems whose solution is easy to check. However, the security of cryptosystems can be studied by means of some security notions. There are three cases which have to be taken into account to study a cryptosystem’s security. In all of them, the adversary who wants to break the security is modeled by a probabilistic polynomial time algorithm. The scheme is secure if the algorithm has no advantage for discovering the secret over a random guesser algorithm, i.e. it is said Eve has a negligible advantage if she guess the correct answer with probability $1/2+\epsilon(k)$, where $k$ is a security parameter and $\epsilon(k)$ is a negligible function. Following this logic, the three security notions considered are:
@@ III-B Quantum Apocalypses
As explained before, the key exchange algorithm for establishing the secret key of a symmetric protocol is done with a public key (asymmetric) cryptosystem. In this sense, doubling the size of the key to be shared increments the complexity of the key generation algorithm of the public key scheme. However, increasing the computational cost is feasible in order to hold the security in almost all the cases and, hence, this is not problematic. Nonetheless, the security of a public key cryptosystem relies on the complexity of the hard problem which has to be solved to get the private key. For example, a brute force attack to RSA consists in trying all primes $p$ and checking whether $p$ is a factor of $N$, requiring $\sqrt{N}$ attempts in the worst case scenario, which is exponential in the digits of $N$ ($d$) and, thus, is an unfeasible task for any classical computer. The most efficient classical algorithm to solve such problem, known as the general number field sieve [63], achieves a complexity of $\mathcal{O}\left(e^{\sqrt{\ln(n)\ln(\ln(n))}}\right)$ asymptotically [64], using this method up to RSA-250 (250 digits) have been factorized satisfactorily [65].
@@
A large enough fault-tolerant quantum computer could run Shor’s algorithm, which is able to factor a number in its primes taking advantage of the laws of quantum mechanics. The first large enough quantum computer will be able to solve factorization problem using $10d$ logical qubits [^6] Note that this number refers to logical or “noiseless” qubits [66]. For quantum computers to work, quantum error correction is required implying that many more physical or “noisy” qubits will be required for an implementation of the Shor algorithm that cracks RSA [67]., where $d$ is the number of digits, with a complexity of $\mathcal{O}(d^{3})$ [64]. A comparison between both algorithms, in terms of their bounds in number of steps, is represented in Fig. 4, where the huge difference (exponential speedup) between both algorithms when the number of digits increases can be easily observed.
Therefore, Shor’s algorithm is a potential threat to the security of all communications over the world, and will be a real problem in the near future since it is able to break most used public key cryptosystems, RSA and ECC schemes. As a result, governments and private companies, such as technological giants IBM, Google and Microsoft, to cite some of them; are dedicating an immense amount of economical and human resources to develop a large enough functional quantum computer (fault-tolerant) able to compute Shor’s algorithm and break the current cybersecurity. Hence, new cryptography schemes are required in order to protect our communications and documents in the post-NISQ [^7] Noisy intermediate-scale quantum era: refers to time scale in which quantum advantage has been firmly proven, but the quantum computers available are still small and too noisy to offer the full potential of quantum computing [68]. NISQ is where we stand right now. quantum era. Another issue to take into account is the fact that some people speculate about the possibility that some global agents are storing encrypted communications for decrypting it in the future once fault-tolerant quantum computers are available. Thus, finding an alternative to RSA and ECC cryptography resistant to quantum attacks is a very relevant problem nowadays. Moreover, it is estimated that a secret holds its value for 15 years [16] and, thus, it is necessary to consider public key cryptosystems based on hard problems that can not be solved efficiently by quantum computers as soon as possible, i.e. it is necessary to implement quantum resistant algorithms 15 years before the first functional quantum computer is available. Google, IBM, China and Xanadu have shown quantum advantage using their computers and they expect a huge increase in their quantum computer capability during the following years [9, 10, 11, 12, 13]. To sum up, we are at a critical moment for cryptography.
@@ IV Post-Quantum Cryptography
Post Quantum Cryptography (PQC) refers to classical cryptographic methods based on hard problems whose solution cannot be found in polynomial time by neither classical nor quantum computers. The hardness of the problem is defined by the computational complexity of an algorithm capable to solve it. In this sense, a quantum computer should not provide any advantage (or such should be almost negligible) in solving the hard problem that stands at the core of a PQC protocol. Note that, as we commented in Section III-B for the case of symmetric cryptography, doubling the key length seems to be enough to keep the security level of classical methods against Grover search attacks. However, CI networks present strong latency requirements implying that doubling the key length could not be an acceptable solution. In this context, there are some proposals for lowering the computational and memory requirements of symmetric key cryptosystems, named Lightweight cryptography (LWC) [70]. LWC reduces the required data to achieve secure communication channels and, thus, reduces the needed device computational resources [71, 72]. Furthermore, it is applicable to networks with legacy and computationally limited devices, including CI networks. Recently, the NIST has finished its standardisation process for LWC [73] including some quantum resilient methods such as Ascon-80pq [74]. In this work we will focus on the key agreement processes (asymmetric key schemes), which stand as more problematic in the context of industrial OT communications due to their latency and computational requirements.
In the following section we will introduce the different hard problems in which PQC methods are based. It is believed that such hard problems are secure against classical and quantum attacks. However, as explained earlier, the security of a cryptosystem is based on some assumption instead of mathematical proofs, hence, it is not possible to assure that any new proposed cryptosystem is secure. Hybrid cryptosystems were proposed in order to keep the current cybersecurity and add quantum resilient protocols, if the PQC protocol is proven to be insecure under classical and quantum attacks in the future the communication protocol would maintain the current security. Hybrid cryptosystems were also conceived to help to the transition from classical to quantum cryptography [75]. Despite of being a very active field and a good proposal to OT cybersecurity, we will not talk about hybrid solutions. Nonetheless, we encourage the interested reader to read the following papers [75, 76, 77].
@@
IV-A Hash-based cryptography
@@
Hash functions are usually employed to create digital signatures, deemed as hash-based signatures (HBS), providing authenticity and integrity to the communication. The first signature scheme using hash-functions was introduced by Lamport in 1979 [80]. Those functions can be classified into One Way Hash Functions (OWHF), Collision Resistant Hash Functions (CRHF) and Universal One way Hash Functions (UOWHF). The general scheme of a hash function cryptosystem is composed by three algorithms, Gen, Enc and Dec in the case of encryption or Gen, Sign and Vry in the case of a digital signature scheme:
@@ IV-A1 Security Notion
IV-A2 PQC protocols
There are two types of hash-based digital signatures (HBS): the One-Time Signature (OTS) scheme and the Multi-Time Signature (MTS) scheme, sometimes called statefull HBS and stateless HBS, respectively. The principal difference between them is the times that a secret key can be used without losing the security assumption. Only Europe, Japan and the United States have a HBS algorithm in their standardization process, SPHINCS+ [82], which is a MTS signature scheme. The parameters of such algorithm is available in Table II. SPHINCS+ has been implemented on an Artix-7 FPGA [83] and on an ARM CortexM3 [84].
IV-B Lattice-based cryptography
@@
Another lattice-based cryptosystem is based on the Learning With Errors (LWE) problem, which was proposed by Regev [87] and which will be later described. The three algorithms involving such cryptography scheme are:
@@ IV-B1 Security Notion
Despite of the existence of an algorithm which finds the non-zero vector of the SVP in time $\mathcal{O}(2^{n})$ [88], there is no quantum algorithm providing an exponential speedup. Indeed, it is worth noting that the SVP problem can be reduced to the CVP problem. In 1996, Ajtai et al. showed that there is a connection between the worst-case problems and the average-case problems in lattice-based cryptography and explain how to construct hard lattice instances from random instances [89]. This connection implies that if there is a probabilistic algorithm that solves the hard problems in the average-case, then there exists a solution for the worst-case scenario. However, it has been proven that there is no probabilistic algorithm in polynomial time for the worst-case scenario. Finally, they conclude that a lattice-based cryptography scheme based on average-case problems can be designed with the security of the worst-case problems. The average-case problems related with lattices are:
@@
IV-B2 PQC protocols
Lattice-based cryptography can be splitted in two diffferent algorithms: NTRU, developed by mathematicians Jeffrey Hoffstein, Jill Pipher, and Joseph H. Silverman in 1996 [86]; and Learning With Errors (LWE) first introduced by Oded Regev in 2009 [87]. Lattice-based cryptography is the most promising quantum resistant cryptosystem, for encryption and signatures. This can be seen as different countries are proposing a large amount of lattice-based protocols to be be standardized. The NTRU encryption protocols that have been proposed to be standardized are: NTRU [91], NTRU-HRSS [92], NTRU-Prime [93] and NTRU+ [94]. The following NTRU protocols have been also proposed to generate signatures: Falcon [95], FatSeal [96], Peregrine [97] and SOLMAE [98]. On the other hand, the LWE encryption protocols proposed to be standardized are: CRYSTALS-Kyber [26], Saber [99], FrodoKEM [100], LAC PKE [101], Aigis-Enc [102] , AKCN-MLWE [103], TALE [104] , AKCN-E8 [105] SCloud [106] and SMAUG-T, which is a merge of SMAUG [107] and [108]; and the following LWE protocols have been also proposed to generate signatures: CRYSTALS-Dilithium [109], Aigis-Sig [110], Mulan [111], NCC-Sign [112] and HAETAE [113].
The following lattice-based PQC protocols have been implemented in HW devices, some of them in devices with low computational resources: NTRU, NTRU-HRSS and NTRU-Prime on an Artix-7 and on aZynq UltraScale+ [114]; NTRU+ on a Xilinx Zynq-7000 [115]; Falcon on an ARM Cortex-A53 [116] and on an ARM CortexM3[84]; CRYSTALS-Kyber on a Xilinx Artix-7 [117], on a Xilinx Artix-7 and on a Virtex-7 FPGAs [118] and on 64-bit ARM Cortex-A processors [119] using number-theoretic transform (NTT) optimization [120, 121], Saber on a Xilinx UltraScale+ [122] and on an Artix-7 and on a Zynq UltraScale+[114] ; LAC on a Xilinx Zynq-7000 [123] and CRYSTALS-Dilithium in [124] on Virtex UltraScale+ and on an ARM Cortex-M4 [124].
The parameters of those algorithms are available in Table I. In [125] they give a good survey explaining different aspects of lattice-based cryptography as their theory, security and performance. They include NIST and Chinese standardisation processes.
IV-C Code-based cryptography
@@ IV-C1 Security Notion
IV-C2 PQC protocols
Two Niederreiter protocols have been proposed to be standardized: BIKE [133], proposed to the NIST standardization process [134]: and PALOMA [135], proposed to the Korea PQC standardization process [136]. In the case of McEllice protocols, NIST process is considering Classical McEllice [137] and Hamming Quasi-Cyclic (HQC)protocols [138] and South Korea is considering REDOG [139].
The parameters of those algorithms are available in Table III.The following code-based PQC protocols have been implemented on a Xilinx Artix-7: BIKE [140], the classic McEliece [141] and HQC [142].
IV-D Multivariate cryptography
@@ IV-D1 Security Notion
IV-D2 PQC protocols
There are two multivariate cryptosystems proposed to be standardized: Rainbow [146] and GeMSS [147].An implementation on an ARM CortexM3 of Rainbow and GeMSS is done in [84]. Both of them have been proposed in the NIST standardization process. However, recently, a new paper has been published claiming that the Rainbow protocol can be broken in just one weekend by means of a laptop [148]. On the other hand, South Korea has proposed a multivariate cryptography protocol called MQ-sign [149].
The parameters of those algorithms are available in Table IV.Rainbow has been implemented in a HW device in [150].
IV-E Isogeny-based cryptography
@@ IV-E1 Security Notion
The best algorithm to solve the SSIP and, hence, that breaks isogeny-based cryptography, is based on meet-in-the-middle attack and its complexity is $\mathcal{O}(p^{1/4})$, which requires $\mathcal{O}(p^{1/4})$ storage capability [157], or, in the case of Supersingular Isogeny Key Encapsulation (SIKE), its security is based on van Oorschot-Wiener (vOW) golden collision finding algorithm [158]. The best quantum algorithm is able to find the secret with $\mathcal{O}(p^{1/6})$, which is not a notorious advantage and does not compromise isogeny-based cryptography [157].
IV-E2 PQC protocols
Supersingular Isogeny Key Encapsulation (SIKE) [159] is an encryption scheme based on Supersingular Isogeny Diffie-Hellman protocol proposed for the NIST standardization process. On the other hand, FIBS is a signature scheme based on the same protocol proposed in the South-Korean standardization process [160]. A compressed alternative of SIKE was proposed in 2016 by Axzarderakhsh et al. in [161], requiring to transmit half of the data. This compressed SIKE was implemented in an ARM Cortex-M4 processor [162] and on a Xilinx Virtex-7 [163], the authors claimed that it is the algorithm which introduces the lowest latency to communications due to its low computational requirements and the extremely compact key sizes. There are also works that try to speed-up the algorithm proposed by L. de Feo [154], such as the work of B. Koziel et al. [164].
The parameters of those algorithms are available in Table V.
IV-F Multiparty Computation protocol and Graph-based cryptography